X & Y jointly gaussian zero mean Groom waste common variance 6² & correlation coefficient pto Find Vanance of Z=x² + X Y
Answer #1

Ret x and y ase roudly greesian zero mean raudan variables common variduce g2 and coordalion codficiends sto. Thuso (x,y) NBV(tirta) - 2MxY & 2 St I (P102 EXY (6 728t, det l2²) 2 {0}(t, + Jh)} 224,66-82) (974 8gt, b46303 (9,²4 286, 227 222² force estte 2 24My A Gazethiffon 21,24,9 {oft + 8 h)} & {2 (2+2 6,3 (9,7728&, k2t 22/02 380 (1, +3628²82 (2+286, d2 + 62 ?)? Pao® , +423 MX Y e 2 = (²+ 286, 22 + 22 2)? alal, 2 {r? ( + gdz)} {2lzt f 2,} (627282, 227 2₂202 2014 ts h) 38²5 (627282, d2 + 22702 t24 M (hirta) - e XY 2 te {o? (byt &f)} te 2 (6,27 28 &tzt 22/0² 24, 22l? {rl + Sh}} {0 (22436, 3² CHOSEB 4,790 {20²6, +Slz)}Then, ho 61, 627 23Mxy E(XY) = 2lgad le,con 2440 ha) -go2 E(44) = ati forto = 204+04 : 364 E (x3y) 24 M 4 1/220, 12 = 0 - at2othen Let z= x² + xy var (2) = vas (8²4xy) var(x2) + Varesy) + 2 Cor (x2xy) = E649-{6x2737 Ebr>y) - (E(xy)]²+ a[ E(+²y) - E()